OpenStudy (anonymous):

Reduce sin 4 x cos 3 x + cos 4 x sin 3 x to a single term. a. cos x b. cos 7 x c. sin 7 x d. sin x

OpenStudy (anonymous):

\[\sin(A)\cos(B) + \cos(A)\sin(B) \implies \color{green}{\sin(A + B)}\]

OpenStudy (anonymous):

in color no less!

OpenStudy (anonymous):

Ha ha ha..

OpenStudy (anonymous):

I thought you are green so I will use green @satellite73

OpenStudy (anonymous):

@IloveCharlie there: \(A = 4x\) and \(B = 3x\).. Can you now use the formula..??

OpenStudy (anonymous):

\[\begin{array}l\color{red}{\normalsize\text{s}}\color{orange}{\normalsize\text{i}}\color{#9c9a2e}{\normalsize\text{n}}\color{green}{\normalsize\text{(}}\color{blue}{\normalsize\text{A}}\color{purple}{\normalsize\text{+}}\color{purple}{\normalsize\text{B}}\color{red}{\normalsize\text{)}}\color{orange}{\normalsize\text{}}\end{array}\]